What is the resistance of a copper wire with a length of 40cm and a cross-sectional area of 0.016mm2 if the resistivity of the material is 1.7x10 -8Ω.m? 


1
Expert's answer
2021-02-24T12:49:12-0500

We can find the resistance of the copper wire from the formula:


"R=\rho\dfrac{l}{A},""R=1.7\cdot10^{-8}\ \Omega\cdot m\cdot\dfrac{0.4\ m}{0.016\cdot10^{-6}\ m^2}=0.425\ \Omega."